Find the x- and y- intercepts: 4x-2y=12

This is how I worked the problem but don't think I did something right.
4x-2y=12 4x-2y=12
4x-2(0)=12 4(0)-2y=12
4x=12 2y=12
x=2 y=-4
(2,0) (0,-4)

Answer by stanbon(75887)

Find the x- and y- intercepts: 4x-2y=12
----------------------------
To find the x-intercept, let y = 0.
You get 4x-0=12
x = 3
x-intercept is (3,0)
----------------------------
To find the y-intercept, let x = 0.
You get 0-2y=12
y = -6
y-intercept is (0,-6)
========================
Cheers,
Stan H.